A new cruise ship line just launched 3 new ships: the Pacific Paradise, The carribean paradise and the MED Paradise. The Car. Paradise has 30 more deluxe staterooms than the Pacific Paradise. The Med. Paradise has 37 fewer deluxe staterooms than twice the number of deluxe staterooms on the pacific paradise. find the number of each deluxe staterooms for each of the ships if the total number of deluxe staterooms for the three ships is 613.
1
Expert's answer
2011-07-28T07:12:23-0400
Denote the number of deluxe staterooms of the Pacific Paradise as X, thus the number of deluxe staterooms of the Carribean Paradise is X + 30, of the Med. Paradise - 2X-37. The total number of the rooms is X + X + 30 + 2X - 37 = 613
4X& -7 = 613
4X = 620
X = 155
Answer: the number of deluxe staterooms on the pacific paradise - is 155, the number of deluxe staterooms on the carribean paradise 155 + 30& = 185, the number of deluxe staterooms on the mediterranean paradise2*155-37 = 273.